# Tensor formalism for predicting synaptic connections with ensemble modeling or optimization

### v2

@@ Abstract
Theoretical neuroscientists often try to understand how the structure of a neural network relates to its function by focusing on structural features that would either follow from optimization or occur consistently across possible implementations. Both optimization theories and ensemble modeling approaches have repeatedly proven their worth, and it would simplify theory building considerably if predictions from both theory types could be derived and tested simultaneously. Here we show how tensor formalism from theoretical physics can be used to unify and solve many optimization and ensemble modeling approaches to predicting synaptic connectivity from neuronal responses. We specifically focus on analyzing the solution space of synaptic weights that allow a threshold-linear neural network to respond in a prescribed way to a limited number of input conditions. For optimization purposes, we compute the synaptic weight vector that minimizes an arbitrary quadratic loss function. For ensemble modeling, we identify synaptic weight features that occur consistently across all solutions bounded by an arbitrary ellipsoid. We derive a common solution to this suite of nonlinear problems by showing how each of them reduces to an equivalent linear problem that can be solved analytically. Although identifying the equivalent linear problem is nontrivial, our tensor formalism provides an elegant geometrical perspective that allows us to solve the problem approximately in an analytical way or exactly using numeric methods. The final algorithm is applicable to a wide range of interesting neuroscience problems, and the associated geometric insights may carry over to other scientific problems that require constrained optimization.
@@ I INTRODUCTION
Quantitatively linking the structure of a biological neural network to its function is a central aim of neuroscience. The network of synaptic connections can be measured directly in small circuits and small brains Bargmann and Marder (2013); Galili et al. (2022); Friedrich and Wanner (2021); Abbott et al. (2020), and one approach is to ask how patterns of synaptic connectivity lead to dynamical patterns of neural activity and behavior Parmelee et al. (2022); Curto and Morrison (2019); Litwin-Kumar and Turaga (2019). However, it remains infeasible to densely reconstruct synaptic connectivity in larger brains, and network reconstruction alone is insufficient to predict the computational and dynamical properties of the biological system Bargmann and Marder (2013); Lappalainen et al. (2023). This paper furthers our efforts Biswas and Fitzgerald (2022) on the inverse problem of predicting synaptic connectivity from the network’s functional responses  Friston et al. (2003); Schneidman et al. (2006); Pillow et al. (2008). A significant appeal of the inverse problem is that various techniques make large-scale recordings possible in many brains Abbott and Svoboda (2020). On the other hand, a major challenge is that many different synaptic connections can often underlie the same functional response properties Huang and Ding (2016); Biswas et al. (2020); Biswas and Fitzgerald (2022). A method that can correctly predict synaptic connectivity would have great theoretical and practical value.
@@
A unified theoretical framework for neural network optimization and ensemble modeling would allow theorists to best capitalize on this experimental progress. Rather than having to toil through disjoint concepts and separate laborious calculations, a unified framework would allow theorists to rapidly make predictions from many different theories. Testing all of these predictions against experimental data would allow the field to rapidly map the domain of applicability of various optimization theories and ensemble modeling techniques Biswas et al. (2020). For instance, linear neural network models are often simple enough to calculate entire solution spaces Atick and Redlich (1990); Baldi and Hornik (1989); Hastie et al. (2009), and it is straightforward to use this solution space to find common structure across solutions Biswas and Fitzgerald (2022), to optimize various quadratic loss functions Fisher et al. (2013); Naumann et al. (2016); Hastie et al. (2009), or to find solutions near empirical constraints Doi et al. (2012). Recent theoretical progress suggests that it may be possible to generalize some of these results to threshold-linear neural network models Biswas and Fitzgerald (2022); Curto et al. (2019); Morrison et al. (2016), which improve the linear model’s biological relevance and computational power while retaining significant analytical tractability. This model class neglects many biological details of real neural networks, but some of these details may not be needed to make accurate predictions Billeh et al. (2020); Biswas et al. (2020); Transtrum and Qiu (2014); Lappalainen et al. (2023); Wanner and Friedrich (2020)
@@
Our mathematical formalism relies on three key insights. First, discerning whether a synaptic weight is non-zero across an elliptically bounded set of weight vectors amounts to finding the smallest ellipse (quadratic cost) that permits the weight to be zero. This maps each ensemble modeling problem onto a related optimization problem. Second, each threshold-linear optimization problem reduces to a linear one once we know which inactive neurons are at threshold in the optimal solution. These neurons impose the same equality constraints that they would in a linear neural network; other inactive neurons are automatically sub-threshold. Third, the tensor formalism introduced by Einstein for relativistic calculations in theoretical physics Misner et al. (1973); Synge and Schild (1978); Wald (2010) facilitates the analysis of threshold-linear optimization problems. In particular, the metric tensor makes it easy to minimize arbitrary quadratic loss functions in non-orthogonal coordinate systems, and writing synaptic weight vectors in a coordinate system that mixes contravariant and covariant basis vectors allows us to assess whether we’ve correctly mapped the threshold-linear problem onto its linear counterpart. The tensor formalism also allows us to approximate the exact solution with an interpretable analytic formula that depends on how neural activity correlates within and across patterns.
@@
We begin in Section II by illustrating different predictions from ensemble modeling and constrained optimization in a simple toy problem. In Section III, we then state the problem we want to solve in its most general form. We re-derive the solution space of  Biswas and Fitzgerald (2022) in the tensor formalism in Section IV, and Section V uses this formalism to derive results for linear neural networks. We expect tensor analysis to be new to some readers, so Appendix A compactly summarizes the most important relations and results with easy-to-reference tables. Furthermore, Appendix B briefly reviews the tensor formalism and metric spaces, including coordinate and basis transformations, the metric tensor, and induced geometries on subspaces. Section VI generalizes the results derived in Section V to the nonlinear case. In Sections V and VI, we progressively solve more complex problems by reducing them to simpler problems. Relatedly, Appendix C shows how our results and formalism generalize as we move from optimizing spherical $L^{2}$ norm to optimizing arbitrary elliptic quadratic functions. Section VII further generalizes our formalism to included additional structural constraints, more complex structual predictions, and predictions about neuronal activity. Appendices D and E facilitate conceptual interpretations of several key results.
@@ II A simple illustrative problem
Finally, one can also use ensemble modeling to make predictions about neural activity measurements. For instance, one can also find a different set of $W$-critical values that allow one to predict whether a given neuron will respond to some arbitrary new pattern or not. In our example, for the new pattern in Fig. 1F (blue arrow), as long as the weights reside within the blue bounding circle, we can predict that $y$ will respond. This is because the circle passes through the intersection of the boundary of the light-blue region and the deep yellow ray of solutions, ensuring that no allowed solutions are in the light-blue region. The projection of the solutions along the new blue pattern vector is therefore always positive.
@@ III General Problem
Our goal in this paper will be to generalize the identification of structure-function links that we described in the simple 2D toy model to arbitrary sized networks, neural patterns, and more general optimization and ensemble modeling problems. In previous work Biswas and Fitzgerald (2022), we’ve shown that once one knows how to make such structure-function links in a feedforward network, it also becomes possible to do so in recurrent networks, at least, in the absence of noise. This continues to hold true even as we generalize our framework. For technical simplicity we will therefore focus on threshold-linear feedforward networks.
@@
and find the weight vector, $\vec{w}_{Q}$, in the solution space that minimizes it (purple dot in Fig. 1E). Setting $Q_{mn}=\delta_{mn}$ and $b_{m}=0$, one recovers the $L^{2}$ norm. However, there may be situations where $y$ may be receiving inputs from different types of neuronal populations so that it may become more sensible to assign the cost of different types of synaptic strengths differently leading to an elliptic optimization function. Some problems demand that the weight vector be near some preferred location other than the origin (e.g., a set of measured weights or the weights that minimizes generalization error), or preferentially close in certain directions than others. A general quadratic function such as Eq. (8) can incorporate these features. Fig. 1E depicts the minimization of such a general quadratic function whose constant surfaces represent off-centered ellipses in the two dimensional weight-space (ellipsoids in a higher dimensional settting). The picture shows the smallest ellipse that contains a solution, the purple dot.
@@ IV Solution Space
This provides two alternate ways to characterize the solution space. Every time we have a constrained pattern, the solution space gets restricted to a hyperplane satisfying $\vec{\varepsilon}^{\,\mu}\cdot\vec{w}=y^{\mu}$, which is perpendicular to the corresponding basis vector, $\vec{\varepsilon}^{\,\mu}$. Thus, the constrained patterns together restrict the solution space to an $({\cal N}-{\cal C})$-dimensional space of the intersection of all the constraining hyperplanes. The null responses, on the other hand, only restrict the weight vectors to reside on the side of the hyperplane $\vec{\varepsilon}^{\,\mu}\cdot\vec{w}=0$ that satisfies the inequality in Eq. (14). While the geometry of the solution space is illuminated by considering projections along the complementary basis vectors $\{\vec{\varepsilon}^{\,\mu}\}$, the weight vectors belonging to the solution space are most readily parametrized in terms of the natural basis $\{\vec{\varepsilon}_{\mu}\}$.
@@ V Structural predictions in a linear theory
In this section our goal will be to slowly build towards a general framework that can make structural predictions based on different principle driven optimization and ensemble modeling approach in a linear theory. In the next section we will generalize our results to threshold-linear transfer functions. Fig. 4 shows the conceptual flow of our derivations.
@@ V.1 Minimum $L^{2}$ norm
Please see Appendix B for more details on how the metric components transform under change of coordinates. The metric provides a way to compute distances between any two points in any coordinate system, and in particular the $L^{2}$ weight norm is just the distance of a point in the weight space from the origin:
@@ V.2.1 From sphere to ellipse
then the above problem can be completely recast in a coordinate invariant language, that of finding the synapse vector with minimum norm, $||\vec{w}||_{Q}$, as defined by the above inner product,
$||\vec{w}||_{Q}^{2}=g_{mn}w^{m}w^{n}\ ,$ (44)
@@
Here the subscript $Q$ in $||\vec{w}||_{Q}^{2}$ signals that this is not the Euclidean norm but rather given by the metric associated with the quadratic function $Q^{2}$, $g_{mn}=Q_{mn}$.
@@
$Q_{\min}^{2}=||\vec{y}||_{Q}=y^{\dot{\mu}}y^{\dot{\nu}}\gamma_{\dot{\mu}\dot{\nu}}\ .$ (54)
@@ V.3 Computing $Q$-critical values from $Q_{\mathtt{\min}}$
Thus as long as $w_{\min}^{p}\neq 0$, determining $Q_{\mathtt{cr},p}$ for the $p^{th}$ synapse in the enemble modeling approach can be converted to an optimization problem of finding $Q_{\mathtt{\min}}$ for a reduced network where the $p^{th}$ input neuron is absent  [^9] Fig. 5 describes this correspondence in the context of the full nonlinear problem.. If $w_{\min}^{p}=0$, then $Q_{\mathtt{\min}}=Q_{\mathtt{cr},p}$, and the synapse can always be eliminated. Or equivalently, one could impose an additional constraining pattern, $x^{\mu={\cal P}+1}_{m}=\delta^{p}{}_{m}$ to which the target response is zero, $y^{{\cal P}+1}=0$:
@@
In Appendix E we have derived an expression of $Q_{\mathtt{cr},p}$ in terms of various correlations involving the pre- and post-synaptic neuronal activities.
@@ VI.1 Combinatorial algorithm
However, there are two issues with this algorithm that one has to address. First, the number of boundaries grows exponentially with ${\cal S}$, which makes this combinatorial algorithm untenable as the network size grows. Secondly, we note that for a given boundary choice, it may turn out that the point where the growing $Q$-surface touches the hyperplane specified by $\eta^{\mu}=0\ \forall\ \mu\in{\cal B}$, is not actually in the solution space, i.e., some of the semi-constrained coordinates that do not belong to ${\cal B}$ may have positive values. We will have to discard these $Q_{{\cal B}}$’s in our combinatorial search algorithm. For e.g. in Fig. 5B (top), as the circular bound grows, it first touches the $\eta^{1s}=0$ line, but the contact point lies outside the solution space as $\eta^{2s}>0$. Thus this contact point has to be ignored.
This problem setup makes it clear that computing $Q_{\mathtt{\min}}$ for a threshold-linear network is equivalent to minimizing $Q$ subject to the equality contraints that $\eta^{\mu}=y^{\mu}$ for all constrained dimensions and the inequality constraints that $\eta^{\mu}\leq 0$ for all semi-constrained dimensions. Since $Q$ is a quadratic function and all of the constraint equations are linear, this is a quadratic programming problem that can be solved numerically with methods from convex optimization Boyd and Vandenberghe (2004); Wright (1999). The combinatorial algorithm explained here is thus wildly inefficient, and it’s akin to a brute force scan through the possible sets of active inequality constraints in the Karush-Kuhn-Tucker conditions Karush (1939); Kuhn and Tucker (1951). Fortunately, the geometric insights that underlie the combinatorial algorithm can be pushed further, and we have discovered a geometric algorithm that can identify the correct boundary via a polynomial search, as well as an analytical formula that often approximates the solution well. We will discuss these algorithms momentarily.

### v3

@@ Abstract
Theoretical neuroscientists often try to understand how the structure of a neural network relates to its function by focusing on structural features that would either follow from optimization or occur consistently across possible implementations. Both optimization theories and ensemble modeling approaches have repeatedly proven their worth, and it would simplify theory building considerably if predictions from both theory types could be derived and tested simultaneously. Here we show how tensor formalism from theoretical physics can be used to unify and solve many optimization and ensemble modeling approaches to predicting synaptic connectivity from neuronal responses. We specifically focus on analyzing the solution space of synaptic weights that allow a threshold-linear neural network to respond in a prescribed way to a limited number of input conditions. For optimization purposes, we compute the synaptic weight vector that minimizes an arbitrary quadratic loss function. For ensemble modeling, we identify synaptic weight features that occur consistently across all solutions bounded by an arbitrary ellipsoid. We derive a common solution to this suite of nonlinear problems by showing how each of them reduces to an equivalent linear problem that can be solved analytically. Although identifying the equivalent linear problem is nontrivial, our tensor formalism provides an elegant geometrical perspective that allows us to solve the problem approximately in an analytical way or exactly using numeric methods. The final algorithm is applicable to a wide range of interesting neuroscience problems, and the associated geometric insights may carry over to other scientific problems that require constrained optimization. We conclude by applying and testing our ensemble modeling framework to whole-brain recordings of larval zebrafish performing optomotor and optokinetic responses.
@@ I INTRODUCTION
Quantitatively linking the structure of a biological neural network to its function is a central aim of neuroscience. The network of synaptic connections can be measured directly in small circuits and small brains Bargmann and Marder (2013); Galili et al. (2022); Friedrich and Wanner (2021); Abbott et al. (2020), and one approach is to ask how patterns of synaptic connectivity lead to dynamical patterns of neural activity and behavior Parmelee et al. (2022); Curto and Morrison (2019); Litwin-Kumar and Turaga (2019). However, it remains infeasible to densely reconstruct synaptic connectivity in larger brains, and network reconstruction alone is insufficient to predict the computational and dynamical properties of the biological system Bargmann and Marder (2013); Lappalainen et al. (2023). This paper furthers our efforts Biswas and Fitzgerald (2022) on the inverse problem of predicting synaptic connectivity from the network’s functional responses  Friston et al. (2003); Schneidman et al. (2006); Pillow et al. (2008). A significant appeal of the inverse problem is that various techniques make large-scale recordings possible in many brains Abbott and Svoboda (2020). A major challenge of this approach is that many different synaptic connections can often underlie the same functional response properties Huang and Ding (2016); Biswas et al. (2020); Biswas and Fitzgerald (2022). Thus a method that can correctly predict synaptic connectivity, or some of their key features, would have great theoretical and practical value.
@@
A unified theoretical framework for neural network optimization and ensemble modeling would allow theorists to best capitalize on this experimental progress. Rather than having to toil through disjoint concepts and separate laborious calculations, a unified framework would allow theorists to rapidly make predictions from many different theories. Testing all of these predictions against experimental data would allow the field to rapidly map the domain of applicability of various optimization theories and ensemble modeling techniques Biswas et al. (2020). For instance, linear neural network models are often simple enough to calculate entire solution spaces Atick and Redlich (1990); Baldi and Hornik (1989); Hastie et al. (2009), and it is straightforward to use this solution space to find common structure across solutions Biswas and Fitzgerald (2022), to optimize various quadratic loss functions Fisher et al. (2013); Naumann et al. (2016); Hastie et al. (2009), or to find solutions near empirical constraints Doi et al. (2012). Recent theoretical progress suggests that it may be possible to generalize some of these results to threshold-linear neural network models Biswas and Fitzgerald (2022); Curto et al. (2019); Morrison et al. (2016), which improve on the linear model’s biological relevance and computational power while retaining significant analytical tractability. This model class neglects many biological details of real neural networks, but some of these details may not be needed to make accurate predictions Billeh et al. (2020); Biswas et al. (2020); Transtrum and Qiu (2014); Lappalainen et al. (2023); Wanner and Friedrich (2020)
@@
Our mathematical formalism relies on three key insights. First, discerning whether a synaptic weight is non-zero across an elliptically bounded set of weight vectors amounts to finding the smallest ellipse (quadratic cost) that permits the weight to be zero. This maps each ensemble modeling problem onto a related optimization problem. This realization also allows us to make new types of predictions that were not discussed in our previous work Biswas and Fitzgerald (2022), such as making predictions about a pattern of synapses (not just a single synaptic connection), or about how a neuron will respond to a new pattern (active or inactive). Such activity predictions allow us to test our ensemble modeling framework without synapse-resolution anatomical data, as illustrated through an application to the neuronal responses underlying the optokinetic Kubo et al. (2014); Svara et al. (2022) and optomotor responses Naumann et al. (2016) of larval zebrafish. Second, each threshold-linear optimization problem reduces to a linear one once we know which inactive neurons are at threshold in the optimal solution. These neurons impose the same equality constraints that they would in a linear neural network; other inactive neurons are automatically sub-threshold. Third, the tensor formalism introduced by Einstein for relativistic calculations in theoretical physics Misner et al. (1973); Synge and Schild (1978); Wald (2010) facilitates the analysis of threshold-linear optimization problems. In particular, the metric tensor makes it easy to minimize arbitrary quadratic loss functions in non-orthogonal coordinate systems, and writing synaptic weight vectors in a coordinate system that mixes contravariant and covariant basis vectors allows us to assess whether we’ve correctly mapped the threshold-linear problem onto its linear counterpart. The tensor formalism also allows us to approximate the exact solution with an interpretable analytic formula that depends on how neural activity correlates within and across patterns.
@@
We begin in Section II by illustrating different predictions from ensemble modeling and constrained optimization in a simple toy problem. In Section III, we then state the problem we want to solve in its most general form. We re-derive the solution space of  Biswas and Fitzgerald (2022) in the tensor formalism in Section IV, and Section V uses this formalism to derive results for linear neural networks. We expect tensor analysis to be new to some readers, so Appendix A compactly summarizes the most important relations and results with easy-to-reference tables. Furthermore, Appendix B briefly reviews the tensor formalism and metric spaces, including coordinate and basis transformations, the metric tensor, and induced geometries on subspaces. Section VI generalizes the results derived in Section V to the nonlinear case. In Sections V and VI, we progressively solve more complex problems by reducing them to simpler problems. Relatedly, Appendix C shows how our results and formalism generalize as we move from optimizing spherical $L^{2}$ norm to optimizing arbitrary elliptic quadratic functions. Section VII further generalizes our formalism to include additional structural constraints, more complex structural predictions, and predictions about neuronal activity. Section VIII illustrates an application of our formalism to neuronal activity data from larval zebrafish performing optokinetic and optomotor behaviors. Finally, Sections IX and X are the Discussion and Methods. Appendices D and E facilitate conceptual interpretations of several key results. Appendix F compares the numerical run times of several algorithms discussed in the paper.
@@ II A simple illustrative problem
Finally, one can also use ensemble modeling to make predictions about neural activity measurements. For instance, one can find a different set of $W$-critical values that allow one to predict whether a given neuron will respond to some arbitrary new pattern or not. In our example, for the new pattern in Fig. 1F (blue arrow), as long as the weights reside within the blue bounding circle, we can predict that $y$ will respond. This is because the circle passes through the intersection of the boundary of the light-blue region and the deep yellow ray of solutions, ensuring that no allowed solutions are in the light-blue region. The projection of the solutions along the new blue pattern vector is therefore always positive.
@@ III General Problem
Our goal in this paper will be to generalize the identification of structure-function links that we described in the simple 2D toy problem to arbitrary sized networks, neural patterns, and more general optimization and ensemble modeling problems. In previous work Biswas and Fitzgerald (2022), we’ve shown that once one knows how to make such structure-function links in a feedforward network, it also becomes possible to do so in recurrent networks, at least, in the absence of noise. This continues to hold true even as we generalize our framework. For technical simplicity we will therefore focus on threshold-linear feedforward networks.
@@
and find the weight vector, $\vec{w}_{Q}$, in the solution space that minimizes it (purple dot in Fig. 1E). Setting $q_{mn}=\delta_{mn}$ and $b_{m}=0$, one recovers the $L^{2}$ norm. However, there may be situations where $y$ is receiving inputs from different types of neuronal populations so that it may become more sensible to assign the cost of different types of synaptic strengths differently leading to an elliptic optimization function. Some problems demand that the weight vector be near some preferred location other than the origin (e.g., a set of measured weights or the weights that minimizes generalization error), or preferentially close in certain directions than others. A general quadratic function such as Eq. (8) can incorporate these features. Fig. 1E depicts the minimization of such a general quadratic function whose constant surfaces represent off-centered ellipses in the two dimensional weight-space (ellipsoids in a higher dimensional settting). The picture shows the smallest ellipse that contains a solution, the purple dot.
@@ IV Solution Space
This provides two alternate ways to characterize the solution space. Every time we have a constrained pattern, the solution space gets restricted to a hyperplane satisfying $\vec{\varepsilon}^{\,\mu}\cdot\vec{w}=y^{\mu}$, which is perpendicular to the corresponding basis vector, $\vec{\varepsilon}^{\,\mu}$. Thus, the constrained patterns together restrict the solution space to an $({\cal N}-{\cal C})$-dimensional space of the intersection of all the constraining hyperplanes. The null responses, on the other hand, only restrict the weight vectors to reside on the side of the hyperplane $\vec{\varepsilon}^{\,\mu}\cdot\vec{w}=0$ that satisfies the inequality in Eq. (14). While the geometry of the solution space is illuminated by considering projections along the complementary basis vectors $\{\vec{\varepsilon}^{\,\mu}\}$, the weight vectors belonging to the solution space are most readily parametrized in terms of the natural basis $\{\vec{\varepsilon}_{\mu}\}$  [^3] The covariant basis vectors also carry biological information. For $\mu=1,\dots,{\cal P}$, the $\mu^{th}$ covariant basis vector is the direction in weight subspace spanned by the patterns that is orthogonal to all the patterns except the $\mu^{th}$ pattern. Therefore, it is the direction for which all other patterns provide no constraints. .
@@ V Structural predictions in a linear theory
In this section our goal will be to slowly build towards a general framework that can make structural predictions based on different principle driven optimization and ensemble modeling approaches in a linear theory. In the next section we will generalize our results to threshold-linear transfer functions. Fig. 4 shows the conceptual flow of our derivations.
@@ V.1 Minimum $L^{2}$ norm
Please see Appendix B for more details on how the metric components transform under change of coordinates. The metric provides a way to compute distances between any two points in any coordinate system, and in particular the $L^{2}$ weight norm is just the Euclidean distance of a point in the weight space from the origin:
@@ V.2.1 From sphere to ellipse
then the above problem can be completely recast in a coordinate invariant language, that of finding the synapse vector with minimum norm, $||\vec{w}||$, as defined by the above inner product,
$||\vec{w}||^{2}=g_{mn}w^{m}w^{n}\ ,$ (44)
@@
$Q_{\min}^{2}=||\vec{y}||_{q}=y^{\dot{\mu}}y^{\dot{\nu}}\gamma_{\dot{\mu}\dot{\nu}}\ ,$ (54)
where the subscript $q$ in $||\vec{y}||_{q}^{2}$ signals that this is not the Euclidean norm but rather given by the metric associated with the quadratic function $Q^{2}$, $g_{mn}=q_{mn}$.
@@ V.3 Computing $Q$-critical values from $Q_{\mathtt{\min}}$
Thus as long as $w_{\min}^{p}\neq 0$, determining $Q_{\mathtt{cr},p}$ for the $p^{th}$ synapse in the ensemble modeling approach can be converted to an optimization problem of finding $Q_{\mathtt{\min}}$ for a reduced network where the $p^{th}$ input neuron is absent Or equivalently, one could impose an additional constraining pattern, $x^{\mu={\cal P}+1}_{m}=\delta^{p}{}_{m}$ to which the target response is zero, $y^{{\cal P}+1}=0$:
@@
In Appendix E we have derived an expression of $Q_{\mathtt{cr},p}$ in terms of various correlations involving the pre- and post-synaptic neuronal activities. Finally, we note that if $w_{\min}^{p}=0$, then $Q_{\mathtt{\min}}=Q_{\mathtt{cr},p}$, and the synapse can always be eliminated.
@@ VI.1 Combinatorial algorithm
However, there are two issues with this algorithm that one has to address. First, the number of boundaries grows exponentially as $2^{{\cal S}}$. Equivalently, there are $2^{{\cal S}}$ ways to assign subsets of semi-constrained dimensions as constrained when converting the threshold-linear optimization problem to a linear optimization problem. This makes the combinatorial algorithm untenable as the network size grows. Secondly, we note that for a given boundary choice, it may turn out that the point where the growing $Q$-surface touches the hyperplane specified by $\eta^{\mu}=0\ \forall\ \mu\in{\cal B}$, is not actually in the solution space, i.e., some of the semi-constrained coordinates that do not belong to ${\cal B}$ may have positive values. We will have to discard these $Q_{{\cal B}}$’s in our combinatorial search algorithm. For e.g. in Fig. 5B (top), as the circular bound grows, it first touches the $\eta^{1s}=0$ line, but the contact point lies outside the solution space as $\eta^{2s}>0$. Thus this contact point has to be ignored.
This problem setup makes it clear that computing $Q_{\mathtt{\min}}$ for a threshold-linear network is equivalent to minimizing $Q$ subject to the equality contraints that $\eta^{\mu}=y^{\mu}$ for all constrained dimensions and the inequality constraints that $\eta^{\mu}\leq 0$ for all semi-constrained dimensions. Since $Q$ is a quadratic function and all of the constraint equations are linear, this is a quadratic programming problem that can be solved numerically with methods from convex optimization Boyd and Vandenberghe (2004); Wright (1999). The combinatorial algorithm explained here is thus wildly inefficient, and it’s akin to a brute force scan through the possible sets of active inequality constraints in the Karush-Kuhn-Tucker conditions Karush (1939); Kuhn and Tucker (1951). However, the geometric insights that underlie the combinatorial algorithm can be pushed further, and we have discovered a geometric algorithm that can identify the correct boundary via a polynomial search, as well as interpretable analytical formulas that often approximates the solution well. We want to emphasize that our goal is not to try to find a numerical algorithm that is more efficient than quadratic programming but rather obtain semi-analytic results that lends biological interpretability in conjunction with a tractable numerical prescription.